1) the ones digit of a two digit number is one more than twice the tens digit. the number is 3 less than four times the sum of the digits of the number what is the number?

2) The ones digit of a two digit number is two times the tens digit. the reverse of the twwo digit number would be twenty seven more than the number. What is the number?

Answer by lwsshak3(11628)

1) the ones digit of a two digit number is one more than twice the tens digit. the number is 3 less than four times the sum of the digits of the number what is the number?
let u=ones digit
let t=tens digit
..
u=2t+1
10t+u=4(t+u)-3
10t+u=4t+4u-3
6t-3u=-3
6t-3(2t+1)=-3
6t-6t-3=-3
-3=-3
no solution
2) The ones digit of a two digit number is two times the tens digit. the reverse of the twwo digit number would be twenty seven more than the number. What is the number?
..
u=2t
10u+t=10t+u+27
9u-9t=27
18t-9t=27
9t=27
t=3
u=2t=6
number: 36